OpenStudy (anonymous):

Which of the following is the slope between the points (2, 3)and (-3, 0)? Answer - 5 over 3 5 over 3 - 3 over 5 3 over 5 -

OpenStudy (anonymous):

last one

OpenStudy (anonymous):

3 over 5

OpenStudy (amistre64):

thats a big question

OpenStudy (anonymous):

\[2,3 and -3,0\]\[0-3\]\[-3-2\] 0-3= -3 -3-2=-5 -3/-5 = 3/5 so the slope is 3 over 5

OpenStudy (anonymous):

thanks